OpenStudy (anonymous):

The perimeter of a rectangular concrete slab is 114 feet and its area is 702 square feet. Find the dimensions of the rectangle. a. Using l for the length of the rectangle, write an expression for the width of the rectangle in terms of l. (Hint: Solve the formula for w.) Show your work. b. Write a quadratic equation using l, the expression you found in part (a), and the area of the slab. c. Solve the quadratic equation. Use the two solutions to find the dimensions of the rectangle.

OpenStudy (anonymous):

a) perimeter 2L + 2w = 114 2w = 114 - 2L w = 57 - L b) for the area Lw = 702 L(57 - L) = 702 57L - L² = 702 L² - 57L + 702 =0 c) quadratic equations L = (57 ±√(-57² - 4(1)(702))) / (2(1)) L = 39 L = 18 the rectangle is 18 by 39 feet we could have factored the equation to solve as well L² - 57L + 702 =0 (L - 18)(L - 39) = 0 L = 18 or L = 39